OpenStudy (anonymous):

what would the roots of 2x^2=-3(2x+3) be in a+b form?

jhonyy9 (jhonyy9):

2x2+6x+9=0 D=36-72=-36 x1,2=(-6+/-sqrt(-36))/4 =(-6+/-6i)/4 =2(3+/-3i)/4 =(3+/-3i)/2

jhonyy9 (jhonyy9):

not is a+b form this is a+bi form